# Improving Implicit Regularization of SGD with Preconditioning for Least Square Problems

### v1

@@ 4 Main Results
In this section, we present the main results of this paper. Detailed proof of the results can be found in the appendix.
Overview of the results. We first present the excessive risk characterization of SGD and ridge with preconditioning. Given our primary objective of demonstrating the superiority of SGD over ridge regression, we focus on establishing an excessive risk upper bound for SGD and an excessive risk lower bound for ridge regression. These bounds provide rigorous support for our analytical framework and claims. Then, building upon the insights gained from the excessive risk analysis, we present and discuss our proposed design for the preconditioning matrix. We then prove that under a theoretical setting (with access to the precise information of the covariance matrix), SGD with our proposed precondition matrix can consistently outperform the standard ridge regression. Moreover, under the same setting, we show that SGD with our proposed preconditioning matrix maintains a consistent advantage over ridge regression when ridge regression employs a representative family of conditioning. Lastly, we address the practical implications of our findings. In a real-world setting, where access to precise covariance matrix information is unavailable, we show that our preconditioning matrix design allows for robust estimation using finite unlabelled data. Remarkably, SGD, with our proposed preconditioning, still maintains a theoretical advantage over ridge regression in this practical scenario.
@@ 4.1 Excessive Risk of Preconditioned SGD and Ridge Regression
Recall that we aim to show that SGD with preconditioning is guaranteed to generalize better than ridge regression. Therefore, we will establish an excessive risk upper bound for SGD and an excessive risk lower bound for ridge regression.
@@
Consider ridge regression with parameter $\lambda>0$ and precondition matrix $\mathbf{M}$. Suppose Assumptions 3.1,  3.2 and  3.3 hold. Let
$\widehat{\mathbf{H}}=\mathbf{H}^{1/2}\mathbf{M}^{-1}\mathbf{H}^{1/2},$
and
$\widehat{\mathbf{w}}^{*}=\mathbf{M}^{1/2}\mathbf{w}^{*}.$
For a constant $\widehat{\lambda}$ depending on $\lambda$, $\widehat{\mathbf{H}}$, $N$ , and $k^{*}=\min\{k:N\widehat{\lambda}_{k}\lesssim\widehat{\lambda}\}$, ridge regression has the following excessive risk lower bound:
$\mathrm{RidgeRisk}\gtrsim\mathrm{RidgeBias}+\mathrm{RidgeVariance},$ (4.1)
$\mathrm{RidgeBias}\eqsim\frac{\widehat{\lambda}^{2}}{N^{2}}\cdot\big{\|}\widehat{\mathbf{w}}^{*}\big{\|}_{\widehat{\mathbf{H}}_{0:k^{*}}^{-1}}^{2}+\|\widehat{\mathbf{w}}^{*}\big{\|}_{\widehat{\mathbf{H}}_{k^{*}:\infty}}^{2},$
$\mathrm{RidgeVariance}\eqsim\sigma^{2}\cdot\bigg{(}\frac{k^{*}}{N}+\frac{N}{\widehat{\lambda}^{2}}\sum_{i>k^{*}}\widehat{\lambda}_{i}^{2}\bigg{)},$
@@
Theorem 4.1 provides a sharp characterization of the lower bound of the excessive risk of the preconditioned ridge. It shows that the excessive risk lower bound of the preconditioned ridge regression depends on the eigenspace of the transformed covariance matrix $\widehat{\mathbf{H}}$ and $\widehat{\mathbf{w}}^{*}$. A couple of things should be noted from the result above. First, if we let $\mathbf{M}=\mathbf{I}$, the result above can be reduced to the lower bound of the excessive risk of standard ridge regression. Therefore, Theorem 4.1 is a more general description of the excessive risk lower bound of ridge regression (Tsigler & Bartlett, 2023). Second, the excessive risk bound is composed of two components, bias and variance. Third, the leading ($i\leq k^{*}$) eigenvectors and the tail ($i>k^{*}$) have different effects on both the bias and variance. Lastly, $k^{*}$ is a constant depending on the problem instance. These observations play an important role in our analysis later.
@@
Let $\mathbf{G}$ be a given preconditioned matrix. Consider SGD update rules with (3) and (3.2). Suppose Assumptions 3.1,  3.2 and  3.3 hold. Let
$\widetilde{\mathbf{H}}=\mathbf{G}^{1/2}\mathbf{H}\mathbf{G}^{1/2}$
and
$\widetilde{\mathbf{w}}^{*}=\mathbf{G}^{-1/2}\mathbf{w}^{*}.$
Suppose the step size $\eta$ satisfies $\eta\leq 1/\mathop{\mathrm{tr}}(\widetilde{\mathbf{H}})$. Then SGD has the following excessive risk upper bound for arbitrary $k_{1},k_{2}\in[d]$:
$\mathrm{SGDRisk}\lesssim\mathrm{SGDBias}+\mathrm{SGDVariance},$ (4.2)
$\mathrm{SGDBias}\eqsim\frac{1}{\eta^{2}N^{2}}\cdot\big{\|}e^{-N\eta\widetilde{\mathbf{H}}}\widetilde{\mathbf{w}}^{*}\big{\|}_{\widetilde{\mathbf{H}}_{0:k_{1}}^{-1}}^{2}+\|\widetilde{\mathbf{w}}^{*}\big{\|}_{\widetilde{\mathbf{H}}_{k_{1}:\infty}}^{2},$
$\mathrm{SGDVariance}\eqsim\big{(}\sigma^{2}+\|\widetilde{\mathbf{w}}^{*}\|_{\widetilde{\mathbf{H}}}\big{)}\cdot\bigg{(}\frac{k_{2}}{N}+N\eta^{2}\sum_{i>k_{2}}\widetilde{\lambda}_{i}^{2}\bigg{)},$
@@ 4.2 Design of The Precondition Matrix
    Theorem 4.1 represents an excessive risk lower bound for ridge and Theorem 4.2 represents an excessive risk upper bound for SGD. Furthermore, as both excessive risk bound can be decomposed into bias and variance, to establish the superiority of SGD over ridge regression, it is sufficient to show:
    $\mathrm{SGDBias}$ $\lesssim\mathrm{RidgeBias},$
    $\mathrm{SGDVariance}$ $\lesssim\mathrm{RidgeVariance}.$
@@
    $N_{\mathrm{sgd}}\geq\Theta(\kappa(N_{\mathrm{ridge}}))N_{\mathrm{ridge}},$
    where,
    $\kappa(N_{\mathrm{ridge}})=\frac{\mathop{\mathrm{tr}}(\mathbf{H})}{N_{\mathrm{ridge}}\lambda_{\min\{N_{\mathrm{ridge}},d\}}},$
@@
where $\beta\geq 0$ is a user-define (tunable) constant. To denote the influence of this parameter, we use ${\mathbf{w}}_{\mathrm{sgd}}(N,\mathbf{G};\beta,\gamma)$ to indicate the tunable parameter $\beta$. It can be observed that as $\beta$ approaches zero, $\mathbf{G}$ converges to the identity matrix $\mathbf{I}$, effectively restoring preconditioned SGD to the standard SGD. Conversely, as we increase $\beta$, the relative effect of $\mathbf{I}$ diminishes for the leading eigenspectrum because $\beta\mathbf{H}$ becomes the dominating factor. Consequently, by tuning $\beta$ appropriately, we can increase and flatten the relative signal strength of the leading eigenspace, leveraging the exponential bias decay from SGD and improving $\kappa(N{\mathrm{ridge}})$.
@@ 4.3 Effectiveness of Preconditioned SGD in Theoretical Setting
The following theorem demonstrates that SGD with our proposed preconditioned matrix can indeed outperform standard ridge regression, i.e., achieving comparable or lower excessive risk.
Theorem 4.3 (Preconditioned-SGD outperforms standard ridge regression).
Consider a preconditioned matrix for SGD as,
$\mathbf{G}=(\beta\mathbf{H}+\mathbf{I})^{-1}.$
Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exists a choice of stepsize $\gamma^{*}$ and a choice of $\beta^{*}$ for SGD such that
$\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\mathbf{G};\beta^{*},\gamma^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{I};\lambda)\big{]}.$
Theorem 4.3 illustrates that SGD with preconditioning can indeed outperform ridge regression in terms of excessive risk, which expands the theoretical understanding of the implicit regularization of SGD and affirmingly supports our central research questions regarding the improvement of SGD’s implicit regularization through preconditioning. More specifically, it demonstrates that our proposed preconditioning matrix design effectively increases the relative strength of the leading eigenspace and improves $\kappa(N_{\mathrm{ridge}})$, thereby closing the existing gap between SGD and ridge regression as observed in previous work (Zou et al., 2021a).
@@
Theorem 4.4 (Preconditioned-SGD outperforms preconditioned ridge regression).
Let $\mathbf{M}$ be the preconditioned matrix for ridge. Consider a preconditioned matrix for SGD as
$\mathbf{G}=\mathbf{M}^{-1/2}(\beta\widehat{\mathbf{H}}+\mathbf{I})^{-1}\mathbf{M}^{-1/2},$
where $\widehat{\mathbf{H}}=\mathbf{H}^{1/2}\mathbf{M}^{-1}\mathbf{H}^{1/2}.$ Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exists a choice of stepsize $\gamma^{*}$ and a choice of $\beta^{*}$ for SGD such that
$\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\mathbf{G};\beta^{*},\gamma^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{M};\lambda)\big{]}.$
Theorem 4.4 demonstrates that, for a representative family of preconditioning matrices used in ridge regression, SGD with our proposed precondition matrix can still outperform ridge regression. Specifically, for a given precondition matrix $\mathbf{M}$ applied to ridge regression, we can find and adjust $\mathbf{G}$ for preconditioned SGD in a way that it surpasses the performance of preconditioned ridge regression. Moreover, by substituting the choice of $\mathbf{G}$ from Theorem 4.3 and 4.4 into $\widetilde{\mathbf{H}}=\mathbf{G}^{1/2}\mathbf{H}\mathbf{G}^{1/2}$, it becomes evident that both methods converge to the same canonical form with respect to $\mathbf{H}$ and $\widehat{\mathbf{H}}$. Consequently, the influence of the design of $\mathbf{G}$ in Theorem 4.4 can be reduced to the case presented in Theorem 4.3, with an updated data covariance matrix $\widehat{\mathbf{H}}$.
@@
Figure 1: Generalization performance comparison among SGD, Ridge regression, PreSGD and PreSGD-Est, where the stepsize $\gamma$, controlling factor $\beta$ and regularization parameter $\lambda$ are fine-tuned to achieve the best performance. The problem dimension is $d=200$ and the variance of model noise is $\sigma^{2}=1$. We consider $6$ combinations of $2$ different covariance matrices and $3$ different ground truth model vectors. The generalization performance (excessive risk) is measured using mean squared error (MSE) on an independent test set of size $1000$ that is generated from the same problem instance. A set of unlabelled data of the same size as the training set is drawn from the same problem instance to estimate $\widetilde{\mathbf{G}}$ in (PreSGD-Est). The plots are averaged over $10$ independent runs.
@@ 4.4 Effectiveness of Preconditioned-SGD with Estimation
In the previous sections, we assumed access to the precise information about the data covariance matrix $\mathbf{H}$. This section addresses the more practical scenarios where obtaining the exact data covariance matrix $\mathbf{H}$ is infeasible. In such cases, we consider situations where we have access to a set of inexpensive unlabeled data, denoted as $\{\widetilde{\mathbf{x}}_{i}\}_{i=1}^{M}$, drawn from the same distribution as the training data. Leveraging this set of unlabeled data, we first estimate $\mathbf{H}$ with the unlabeled data and subsequently design the preconditioned matrix using the estimated covariance information. We denote the estimated covariance matrix as $\mathbf{\Sigma}$, defined as follows:
$\mathbf{\Sigma}=M^{-1}\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\top}$
where $\widetilde{\mathbf{X}}=[\widetilde{\mathbf{x}}_{1},...,\widetilde{\mathbf{x}}_{M}]\in\mathbb{R}^{d\times M}$ is a set of $M$ unlabeled data, which is often accessible and cost-effective to obtain in practice. Moreover, in the case where these data are unavailable, we can use half of the training data to approximate $\bm{\Sigma}$ to decouple the randomness in theory. In practice, if using training data to perform both stochastic gradient updates and precondition estimation, the preconditioned SGD would resemble the well-known Newton methods.
@@
The following results demonstrate that our proposed precondition matrix allows for robust estimation from unlabeled data and still allows preconditioned SGD to maintain a theoretical advantage over ridge regression.
@@
Theorem 4.5 (Preconditioned-SGD with estimated $\widetilde{\mathbf{G}}$ outperforms standard ridge regression).
Consider a preconditioned matrix for SGD as
$\widetilde{\mathbf{G}}=(\beta\mathbf{\Sigma}+\mathbf{I})^{-1}.$
@@
Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exist a choice of stepsize $\gamma^{*}$ and a choice of $\beta^{*}$ for SGD such that with probability at least $1-\exp(-\Omega(M))$, we have,
$\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\widetilde{\mathbf{G}};\beta^{*},\gamma^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{I};\lambda)\big{]}.$
@@
Theorem 4.6 (Preconditioned-SGD with estimated $\widetilde{\mathbf{G}}$ outperforms preconditioned ridge regression).
Let $\mathbf{M}$ be the preconditioned matrix for ridge. Consider a preconditioned matrix for SGD as
$\widetilde{\mathbf{G}}=\mathbf{M}^{-1/2}(\beta\widehat{\mathbf{H}}+\mathbf{I})^{-1}\mathbf{M}^{-1/2},$
@@
where $\widehat{\mathbf{H}}=\mathbf{\Sigma}^{1/2}\mathbf{M}^{-1}\mathbf{\Sigma}^{1/2}.$ Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exists a choice of stepsize $\gamma^{*}$ and a choice of $\beta^{*}$ for SGD such that with probability at least $1-\exp(-\Omega(M))$,
$\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\widetilde{\mathbf{G}};\beta^{*},\gamma^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{M};\lambda)\big{]}.$
@@
Theorems 4.5 and 4.6 are variations of Theorems 4.3 and  4.4 that incorporate empirical estimation. They establish the efficacy of our proposed preconditioned matrix design and demonstrate that even when the data covariance matrix $\mathbf{H}$ is estimated from unlabeled data, SGD with our proposed preconditioned matrix can still outperform (standard & preconditioned) ridge regression with a theoretical guarantee. The introduction of randomness due to empirical estimation is reflected in the probabilistic bound related to the estimation sample size $M$ in the theorem statements. Nonetheless, despite the similarities in their formal expressions, establishing Theorems4.5 and 4.6 as opposed to Theorems 4.3 and  4.4 presents distinctive technical challenges. Our analytical approach for comparing SGD and ridge relies on the characterization of excessive risk within the eigenspace. The randomness induced by the empirical estimation can introduce stochastic fluctuations to the eigenvalues within the eigenspace, thereby complicating the analysis. Therefore, to properly analyze and compare SGD and ridge under conditions involving empirical estimation, it is imperative to employ a specialized set of technical tools designed to mitigate the impact of the randomness introduced through empirical estimation.

### v3

@@ 4 Main Results
Overview of the results. We first present the excessive risk characterization of SGD and ridge with preconditioning. Given our primary objective of demonstrating the comparable performance of SGD over ridge regression, we focus on establishing an excessive risk upper bound for SGD and an excessive risk lower bound for ridge regression. These bounds provide rigorous support for our analysis and claims. Then, building on the insights from the excessive risk analysis, we discuss and present our design for the preconditioning matrix. We then prove that under a theoretical setting (with access to the precise information of the covariance matrix), SGD with our proposed precondition matrix can consistently achieve comparable performance to standard ridge regression. Moreover, under the same setting, we show that SGD with our proposed preconditioning matrix maintains a comparable performance to ridge regression when ridge regression employs a representative family of conditioning. Lastly, we address the practical implications of our findings. In a real-world setting, where access to precise covariance matrix information is unavailable, we show that our preconditioning matrix design allows for robust estimation using finite unlabelled data. Remarkably, SGD, with our proposed preconditioning, still guarantee to maintain a comparable performance over ridge regression in this practical scenario. Detailed proof can be found in the appendix.
@@ 4.1 Excessive Risk of Preconditioned SGD and Ridge Regression
Consider ridge regression with parameter $\lambda>0$ and precondition matrix $\mathbf{M}$. Suppose Assumptions 3.1,  3.2 and  3.3 hold. Let $\widehat{\mathbf{H}}=\mathbf{H}^{1/2}\mathbf{M}^{-1}\mathbf{H}^{1/2},\quad\text{and}\quad\widehat{\mathbf{w}}^{*}=\mathbf{M}^{1/2}\mathbf{w}^{*}.$ For a constant $\widehat{\lambda}$ depending on $\lambda$, $\widehat{\mathbf{H}}$, $N$ , and $k^{*}=\min\{k:N\widehat{\lambda}_{k}\lesssim\widehat{\lambda}\}$, ridge regression has the following excessive risk lower bound:
$\mathrm{RidgeRisk}\gtrsim\underbrace{\frac{\widehat{\lambda}^{2}}{N^{2}}\cdot\big{\|}\widehat{\mathbf{w}}^{*}\big{\|}_{\widehat{\mathbf{H}}_{0:k^{*}}^{-1}}^{2}+\|\widehat{\mathbf{w}}^{*}\big{\|}_{\widehat{\mathbf{H}}_{k^{*}:\infty}}^{2}}_{\mathrm{RidgeBias}}+\underbrace{\sigma^{2}\cdot\bigg{(}\frac{k^{*}}{N}+\frac{N}{\widehat{\lambda}^{2}}\sum_{i>k^{*}}\widehat{\lambda}_{i}^{2}\bigg{)}}_{\mathrm{RidgeVariance}},$ (4.1)
@@
Theorem 4.1 provides a sharp characterization of the excessive risk lower bound for the preconditioned ridge. It shows that the excessive risk lower bound of the preconditioned ridge regression depends on the eigenspace of the transformed covariance matrix $\widehat{\mathbf{H}}$ and $\widehat{\mathbf{w}}^{*}$. A couple of things should be noted from the result above. First, if we let $\mathbf{M}=\mathbf{I}$, the result above can be reduced to the lower bound of the excessive risk of standard ridge regression. Therefore, Theorem 4.1 is a more general description of the excessive risk lower bound of ridge regression (Tsigler & Bartlett, 2023). Second, the excessive risk bound is composed of two components, bias and variance. Third, the leading ($i\leq k^{*}$) eigenvectors and the tail ($i>k^{*}$) have different effects on both the bias and variance. Lastly, $k^{*}$ is a constant depending on the problem instance. These observations play an important role in our analysis later.
@@
Let $\mathbf{G}$ be a given preconditioned matrix. Consider SGD update rules with (3.1) and (3.2). Suppose Assumptions 3.1,  3.2 and  3.3 hold. Let $\widetilde{\mathbf{H}}=\mathbf{G}^{1/2}\mathbf{H}\mathbf{G}^{1/2},\quad\text{and}\quad\widetilde{\mathbf{w}}^{*}=\mathbf{G}^{-1/2}\mathbf{w}^{*}.$ Suppose the step size $\eta$ satisfies $\eta\leq 1/\mathop{\mathrm{tr}}(\widetilde{\mathbf{H}})$. Then SGD has the following excessive risk upper bound for arbitrary $k_{1},k_{2}\in[d]$:
$\mathrm{SGDRisk}\lesssim\underbrace{\frac{1}{\eta^{2}N^{2}}\cdot\big{\|}e^{-N\eta\widetilde{\mathbf{H}}}\widetilde{\mathbf{w}}^{*}\big{\|}_{\widetilde{\mathbf{H}}_{0:k_{1}}^{-1}}^{2}+\|\widetilde{\mathbf{w}}^{*}\big{\|}_{\widetilde{\mathbf{H}}_{k_{1}:\infty}}^{2}}_{\mathrm{SGDBias}}+\underbrace{\big{(}\sigma^{2}+\|\widetilde{\mathbf{w}}^{*}\|_{\widetilde{\mathbf{H}}}\big{)}\cdot\bigg{(}\frac{k_{2}}{N}+N\eta^{2}\sum_{i>k_{2}}\widetilde{\lambda}_{i}^{2}\bigg{)}}_{\mathrm{SGDVariance}},$ (4.2)
@@ 4.2 Design of The Precondition Matrix
    Theorem 4.1 represents an excessive risk lower bound for ridge and Theorem 4.2 represents an excessive risk upper bound for SGD. Furthermore, as both excessive risk bound can be decomposed into bias and variance, to prove that SGD is comparable with ridge regression, it is sufficient to show:
    $\mathrm{SGDBias}\lesssim\mathrm{RidgeBias},\quad\text{and}\quad\mathrm{SGDVariance}\lesssim\mathrm{RidgeVariance}.$
@@
    $N_{\mathrm{sgd}}\geq\Theta(\kappa(N_{\mathrm{ridge}}))N_{\mathrm{ridge}},\quad\text{where}\quad\kappa(N_{\mathrm{ridge}})=\frac{\mathop{\mathrm{tr}}(\mathbf{H})}{N_{\mathrm{ridge}}\lambda_{\min\{N_{\mathrm{ridge}},d\}}},$
@@
where $\beta\geq 0$ is a user-define (tunable) constant. To denote the influence of this parameter, we use ${\mathbf{w}}_{\mathrm{sgd}}(N,\mathbf{G};\beta,\eta)$ to indicate the tunable parameter $\beta$. It can be observed that as $\beta$ approaches zero, $\mathbf{G}$ converges to the identity matrix $\mathbf{I}$, effectively restoring preconditioned SGD to the standard SGD. Conversely, as we increase $\beta$, the relative effect of $\mathbf{I}$ diminishes for the leading eigenspectrum because $\beta\mathbf{H}$ becomes the dominating factor. Consequently, by tuning $\beta$ appropriately, we can increase and flatten the relative signal strength of the leading eigenspace, leveraging the exponential bias decay from SGD and improving $\kappa(N{\mathrm{ridge}})$.
@@ 4.3 Effectiveness of Preconditioned SGD in Theoretical Setting
The following theorem demonstrates that SGD with our proposed preconditioned matrix can indeed be comparable to standard ridge regression, i.e., achieving comparable or lower excessive risk.
Theorem 4.3 (Preconditioned-SGD comparable to standard ridge regression).
Consider a preconditioned matrix for SGD as, $\mathbf{G}=(\beta\mathbf{H}+\mathbf{I})^{-1}.$ Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exists a choice of stepsize $\eta^{*}$ and a choice of $\beta^{*}$ for SGD such that $\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\mathbf{G};\beta^{*},\eta^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{I};\lambda)\big{]}.$
Theorem 4.3 illustrates that preconditioned SGD can indeed be comparable to ridge regression, which expands the theoretical understanding of the implicit regularization of SGD and affirmingly supports our central research questions regarding the improvement of SGD’s implicit regularization through preconditioning. It demonstrates that our proposed preconditioning matrix design effectively increases the relative strength of the leading eigenspace and improves $\kappa(N_{\mathrm{ridge}})$, thereby closing the existing gap between SGD and ridge regression as observed in previous work (Zou et al., 2021a).
@@
Theorem 4.4 (Preconditioned-SGD comparable to preconditioned ridge regression).
Let $\mathbf{M}$ be the preconditioned matrix for ridge. Consider a preconditioned matrix for SGD as $\mathbf{G}=\mathbf{M}^{-1/2}(\beta\widehat{\mathbf{H}}+\mathbf{I})^{-1}\mathbf{M}^{-1/2},$ where $\widehat{\mathbf{H}}=\mathbf{H}^{1/2}\mathbf{M}^{-1}\mathbf{H}^{1/2}.$ Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exists a choice of stepsize $\eta^{*}$ and a choice of $\beta^{*}$ for SGD such that $\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\mathbf{G};\beta^{*},\eta^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{M};\lambda)\big{]}.$
Theorem 4.4 demonstrates that, for a representative family of preconditioning matrices used in ridge regression, SGD with our proposed precondition matrix can still achieve comparable performance with ridge regression. Specifically, for a given precondition matrix $\mathbf{M}$ applied to ridge regression, we can find and adjust $\mathbf{G}$ for preconditioned SGD in a way that matches the performance of preconditioned ridge regression. Moreover, by substituting the choice of $\mathbf{G}$ from Theorem 4.3 and 4.4 into $\widetilde{\mathbf{H}}=\mathbf{G}^{1/2}\mathbf{H}\mathbf{G}^{1/2}$, it becomes evident that both methods converge to the same canonical form with respect to $\mathbf{H}$ and $\widehat{\mathbf{H}}$. Consequently, the influence of the design of $\mathbf{G}$ in Theorem 4.4 can be reduced to the case presented in Theorem 4.3, with an updated data covariance matrix $\widehat{\mathbf{H}}$. Additionally, one might naturally question whether the reverse direction holds true with preconditioning, i.e., whether ridge regression can similarly be guaranteed to match the performance of preconditioned SGD through appropriate preconditioning and tuning. We demonstrate in Appendix H (Theorem H.1) that this is indeed the case, suggesting an inherent correlation between SGD and ridge regression.
@@
Figure 1: Generalization performance comparison among SGD, Ridge regression, PreSGD and PreSGD-Est, where the stepsize $\eta$, controlling factor $\beta$ and regularization parameter $\lambda$ are fine-tuned to achieve the best performance. The problem dimension is $d=200$ and the variance of model noise is $\sigma^{2}=1$. We consider $6$ combinations of $2$ different covariance matrices and $3$ different ground truth model vectors. The generalization performance (excessive risk) is measured using mean squared error (MSE) on an independent test set of size $1000$ that is generated from the same problem instance. A set of unlabelled data of the same size as the training set is drawn from the same problem instance to estimate $\widetilde{\mathbf{G}}$ in (PreSGD-Est). The plots are averaged over $10$ independent runs.
@@ 4.4 Effectiveness of Preconditioned-SGD with Estimation
In the previous sections, we assumed access to the precise information about the data covariance matrix $\mathbf{H}$. This section addresses the more practical scenarios where obtaining the exact data covariance matrix $\mathbf{H}$ is infeasible. In such cases, we consider situations where we have access to a set of inexpensive unlabeled data, denoted as $\{\widetilde{\mathbf{x}}_{i}\}_{i=1}^{M}$, drawn from the same distribution as the training data. Leveraging this set of unlabeled data, we first estimate $\mathbf{H}$ with the unlabeled data and subsequently design the preconditioned matrix using the estimated covariance information. We denote the estimated covariance matrix as $\mathbf{\Sigma}=M^{-1}\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\top}$, where $\widetilde{\mathbf{X}}=[\widetilde{\mathbf{x}}_{1},...,\widetilde{\mathbf{x}}_{M}]\in\mathbb{R}^{d\times M}$ is a set of $M$ unlabeled data, which is often accessible and cost-effective to obtain in practice. Moreover, in the case where these data are unavailable, we can use half of the training data to approximate $\bm{\Sigma}$ to decouple the randomness in theory. In practice, if using training data to perform both stochastic gradient updates and precondition estimation, the preconditioned SGD would resemble the well-known Newton methods to some extent 333we provide a more in-depth comparison between our method and Newton’s method in Appendix G.
@@
The following results demonstrate that our proposed precondition matrix allows for robust estimation from unlabeled data and still allows preconditioned SGD to maintain a theoretical improvement and remain comparable to ridge regression.
@@
Theorem 4.5 (Preconditioned-SGD with estimated $\widetilde{\mathbf{G}}$ comparable to standard ridge regression).
@@
Consider a preconditioned matrix for SGD as $\widetilde{\mathbf{G}}=(\beta\mathbf{\Sigma}+\mathbf{I})^{-1}.$ Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exist a choice of stepsize $\eta^{*}$ and a choice of $\beta^{*}$ for SGD such that with probability at least $1-\exp(-\Omega(M))$, we have $\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\widetilde{\mathbf{G}};\beta^{*},\eta^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{I};\lambda)\big{]}.$
@@
Theorem 4.6 (Preconditioned-SGD with estimated $\widetilde{\mathbf{G}}$ comparable to preconditioned ridge regression).
@@
Let $\mathbf{M}$ be the preconditioned matrix for ridge. Consider a preconditioned matrix for SGD as $\widetilde{\mathbf{G}}=\mathbf{M}^{-1/2}(\beta\widehat{\mathbf{H}}+\mathbf{I})^{-1}\mathbf{M}^{-1/2},$ where $\widehat{\mathbf{H}}=\mathbf{\Sigma}^{1/2}\mathbf{M}^{-1}\mathbf{\Sigma}^{1/2}.$ Let $N$ be a sufficiently large training sample size for both ridge and SGD. Then, for any ridge regression solution that is generalizable and any $\lambda>0$, there exists a choice of stepsize $\eta^{*}$ and a choice of $\beta^{*}$ for SGD such that with probability at least $1-\exp(-\Omega(M))$, we have $\mathcal{E}\big{[}{\mathbf{w}}_{\mathrm{sgd}}(N,\widetilde{\mathbf{G}};\beta^{*},\eta^{*})\big{]}\lesssim\mathcal{E}\big{[}\mathbf{w}_{\mathrm{ridge}}(N,\mathbf{M};\lambda)\big{]}.$
@@
Theorems 4.5 and 4.6 are variations of Theorems 4.3 and  4.4 that incorporate empirical estimation. They establish the efficacy of our proposed preconditioned matrix design and demonstrate that even when the data covariance matrix $\mathbf{H}$ is estimated from unlabeled data, SGD with our proposed preconditioned matrix can still be comparable to (standard & preconditioned) ridge regression with a theoretical guarantee. The introduction of randomness due to empirical estimation is reflected in the probabilistic bound related to the estimation sample size $M$ in the theorem statements. Nonetheless, despite the similarities in their formal expressions, establishing Theorems 4.5 and 4.6 as opposed to Theorems 4.3 and  4.4 presents distinctive technical challenges. Our analytical approach for comparing SGD and ridge relies on the characterization of excessive risk within the eigenspace. The randomness induced by the empirical estimation can introduce stochastic fluctuations to the eigenvalues within the eigenspace, thereby complicating the analysis. Therefore, to properly analyze and compare SGD and ridge under conditions involving empirical estimation, it is imperative to employ a specialized set of technical tools designed to mitigate the impact of the randomness introduced through empirical estimation.